OpenStudy (anonymous):

A sequence has a second term of 1 and a common difference of -3. The fifth term of the sequence is _____. -11 -8 -5 10

OpenStudy (anonymous):

a+d=1 where a=1st term and d= common difference so a-3=1 soa=4 fifth term will be a+4d=4-12=-8